OpenStudy (anonymous):

Express in trigonometric form. -5i

OpenStudy (anonymous):

a+ib=-5i on comparing a=0 and b=5

OpenStudy (anonymous):

trigo form =r*(cos(theta)+i*sin(theta))

OpenStudy (anonymous):

where \[r=\sqrt{a^{2}+b^{2}}\]

OpenStudy (anonymous):

\[\theta=\tan^{-1} (\frac{ b}{ a })=\tan^{-1} \frac{ 5 }{ 0 }=\tan^{-1} \infty=\frac{ \pi }{ 2}\]

OpenStudy (anonymous):

ab b is negative so actual theta =pi-theta=pi/2

OpenStudy (anonymous):

a+ib=r*(cos(theta)+i*sin(theta)) a=r*cos(theta) b=r*sin(theta)